OpenStudy (anonymous):

Why does raising something to 0 get you one?

OpenStudy (zarkon):

suppose \(a\ne 0\) then \[a^0=a^{1-1}=a^1a^{-1}=\frac{a}{a}=1\]

OpenStudy (anonymous):

wow that's perfect Thanks!